We consider following yield curve

Maturity 1 2 3 4 5 ... 10 .. 50
Yield (%) 12 11,5 11 10,5 10 ... 9,5 ... 9

we are planning to participate in an investment which will be made in 3 years from now.This investment will bring you every year 2 million USD over 7 years. What is the maximum price you will be willing to accept.

Please . Thanks
Answer #1

Interest rate of the proposed investment in three years is the spot rate for 7 years after 3 years from now ie., 3 years forward rate for 7 year maturity (f3,7)

Yield curve depicts the spot rates for various maturities, as of now.

Given, spot rate for 3 years maturity S3= 11% and Spot rate for 10 years S10= 9.5%

Therefore, forwards rate for 7 years after 3 years=

{[(1+S10)^10 / (1+S3)^3]^1/7}-1

(1+9.5%)^10=2.478228

(1+11%)^3=1.367631

(1+9.5%)^10/(1+11%)^3= 2.478228/1.367631= 1.81205867

[(1+9.5%)^10/(1+11%)^3]^1/7=1.812059^(1/7) = 1.088633641

{[(1+9.5%)^10/(1+11%)^3]^(1/7)}-1= 1.088634 – 1 = 0.88633641 Or, 8.86333641%

Using the above rate as interest rate for 7 years, price of the bond is $ 10.11223056 Million calculated as the present value of annuity as follows:

Present Value of Annuity Payments at the end of each period 1 3 Present value of annuity is calculated using the formula PVA=